OpenStudy (anonymous):

The height of a cone is two times its base diameter. What is the volume of the cone in terms of its base radius r?

OpenStudy (mathstudent55):

Do you know the formula of the volume of a cone?

OpenStudy (anonymous):

what are the choices?

OpenStudy (anonymous):

v = π r 2 h 3

OpenStudy (mathstudent55):

is there a fraction there? Is there an exponent there?

OpenStudy (anonymous):

A. V= 2/3 PIE R2 B. V= 4/3 PIE R3 C. V= 1/3 PIE R3 D. V= 1/3 PIE R2

OpenStudy (mathstudent55):

Do you mean this for the formula? \(\LARGE V_{cone} = \dfrac{\pi r^2 h}{3}\)

OpenStudy (anonymous):

YEAH

OpenStudy (mathstudent55):

Great. That is the correct formula. Notice the formula uses r to mean radius of the base, and h to mean the height of the cone.

OpenStudy (mathstudent55):

Now let's look at your problem's cone. The radius is r. What is the diameter, d, compared to the radius?

OpenStudy (mathstudent55):

In any circle, what is the relationship between the diameter and the radius?

OpenStudy (mathstudent55):